OpenStudy (anonymous):

Show that cos(A-B)-cos(A+B)=2sinAsinB

OpenStudy (mertsj):

use these: \[\cos (A-B)=\cos A \cos B+\sin A \sin B\] \[\cos (A+B)=\cos A \cos B-\sin A \sin B\]